5.6 L of helium gas at  STP is adiabatically compressed to 0.7 L.


Taking the initial temp. to be t1 work done in the process is________

Grade:10

4 Answers

Chetan Mandayam Nayakar
12 years ago

let P1,V1,T1 be initial values and P2,V2,T2 be final values. T1*(V1)^(γ-1)=T2*(V2)^(γ-1), γ=5/3,P1*((V1)^γ)=P2*((V2)^γ)

work done= (P2V2-P1V1)/(γ-1)

 

ABHISHEK GAKHAR
12 years ago

give answer as a function of t1

Suraj Negi
12 years ago

Answer : 9/8RT1

HP
8 years ago
hint?